# Variational Prior Replacement in Bayesian Inference and Inversion Xuebin Zhao1 and Andrew Curtis1

### v1

@@ 2.4 Variational Prior Replacement
Solving the second (quasi) Bayesian problem in equation 5 requires to be evaluated. While in most cases both old and new prior probability values can be calculated efficiently, the main challenge lies in computing without invoking equation 4 (Bayes’ rule) which involves forward simulation of a given model sample. Otherwise prior replacement would reduce to solving multiple independent Bayesian inverse problems with different prior distributions.
If we use variational inference to solve the first Bayesian problem involved in prior replacement (equation 4), then
Substitute this into equation 5
@@
Since forward simulation is the most computationally expensive component in an inverse problem, evaluating would normally be much cheaper than evaluating . Equation 11 indicates that we can alternatively calculate from , up to a normalisation constant which can be absorbed into .
@@
Note that not all variational inference methods can provide a variational distribution whose probability value can be evaluated easily. For example, Stein variational gradient descent (SVGD – Liu & Wang, 2016) and its stochastic version (sSVGD – Gallego & Insua, 2018) iteratively update a set of samples (also called particles) such that they become distributed according to a good approximation to the posterior distribution. The output is the optimised set of particles, which are used to estimate statistical information of the posterior distribution. However, evaluating the probability value for a particular is not at all straightforward with these methods.
@@
Another suite of variational methods assumes that the posterior distribution has a known structure and complexity (for example a Gaussian distribution), and can thus be expressed by a given parametric (often analytic) representation. Variational inference finds the optimal values of hyperparameters that control the parametric expression, thus defining a variational distribution that best approximates the true posterior pdf. Since we obtain a parametric (closed form) expression for the variational distribution, we can easily evaluate its probability value for any given sample . We refer to this kind of variational method as parametric variational inference (Sjölund, 2023). Examples of typical parametric variational inference methods include automatic differentiation variational inference (ADVI – Kucukelbir et al., 2017), normalising flows (Rezende & Mohamed, 2015), boosting variational inference (BVI – Guo et al., 2016; Miller et al., 2017), and physically structured variational inference (PSVI – Zhao & Curtis, 2024b).
@@
If is constructed using a parametric variational inference method, the problem described in equation 11 can in principle be solved using any probabilistic inference method. However, even though can be evaluated efficiently, the curse of dimensionality (Curtis & Lomax, 2001) may nevertheless make the problem expensive, if not impossible, to solve using Monte Carlo sampling methods. We therefore introduce a second variational distribution to approximate the new posterior distribution given the new prior information . As introduced in the previous section, this new variational distribution can be obtained by minimising the KL-divergence between and :
@@
where the second line is obtained by substituting equation 11 into the first line. The last term is a constant and can safely be ignored when minimising KL. This optimisation problem can be solved in exactly the same way as conventional variational problems. The optimisation result satisfies
@@
Using the framework of variational inference, the two main operations in the original prior replacement problem described in equations 4 and 5 are converted into two variational problems represented by equations 10 and 13. We therefore call this new methodology variational prior replacement (VPR).
To solve a VPR problem, the first operation (equation 10) which estimates must be performed using a parametric variational method since its probability value needs to be evaluated in the second step (equation 13). The second problem can be solved using any variational method.
@@
The most expensive step in the VPR algorithm is to solve the first variational problem, which requires the likelihood term (forward function) to be calculated. Provided that the first problem is solved with the prior pdf that has the broadest support, and that the support of all other prior pdf’s are subsets of this support, then this step needs to be performed only once, after which we can replace prior information rapidly whenever it changes, without solving a new Bayesian problem again. This makes the proposed method attractive in real problems, especially when multiple different prior distributions are possible for a single observed dataset (Earp & Curtis, 2020; Bloem et al., 2024).
@@ 2.5 Physically Structured Variational Inference (PSVI)
In this paper, we use physically structured variational inference (PSVI – Zhao & Curtis, 2024b) to solve the two variational problems in equations 10 and 13. PSVI is an efficient parametric variational inference method that defines a Gaussian variational family with a physics-based correlation structure. When the model parameters to be estimated have physical constrains (for example, seismic velocity should be a positive number and earthquake source location should be below the Earth’s surface), a bijective function (an invertible transform) is usually applied to the Gaussian random variables to ensure that the transformed model parameters satisfy their physical constrains. For example, the following logit functions
@@
A Gaussian variational distribution is typically defined by a mean vector and a covariance matrix . To ensure that always remains positive semi-definite, we re-parametrise it using a Cholesky factorisation , where is a lower triangular matrix. A full covariance matrix can be constructed to include correlation information between pairs of model parameters. However, this incurs huge memory requirements and computational costs (for an dimensional problem, requires real-valued entries). Alternatively, a mean-field (factorised) Gaussian variational approximation may be used for high dimensional problems, which defines a diagonal covariance matrix, thus ignoring all correlations between model parameters. These two options are referred to as mean-field ADVI and full rank ADVI in Kucukelbir et al. (2017). Unfortunately for full waveform inversion (FWI) problems considered in this paper, mean-field ADVI normally underestimates uncertainties of the posterior distribution whereas full rank ADVI is intractable considering the dimensionality of the problem (Zhang et al., 2023; Zhao & Curtis, 2024b).
PSVI balances mean-field ADVI and full rank ADVI by modelling only the most important (dominant) correlation information in model vector , guided by physical properties (prior knowledge) of imaging problems. Specifically, in spatial inverse (imaging) problems, model correlations are shown to be strong mainly between pairs of locations that are in spatial proximity to each other, and the magnitude of correlations decreases rapidly as the distance between two locations increases (Gebraad et al., 2020; Zhang & Curtis, 2021a; Biswas & Sen, 2022). This suggests that it might be sufficient to model correlations only between parameters that are close (e.g., for FWI, parameters of cells that lie within a dominant wavelength of one another), and ignore correlations between those that are further apart.
@@
For each element, the first subscript indicates a block of off-diagonal elements that are rows below the main diagonal (i.e., at an offset of from the main diagonal), and the second subscript indicates that is the th element of that off-diagonal block. In equation 16, sparsely distributed off-diagonal elements in red are used to capture main correlations between parameter pairs that are assumed to be important (in this case, are spatially close), and these are optimised during variational inference. All other off-diagonal elements of are imposed to be zero by assuming independence of the corresponding model parameter pairs. Note that we only impose a sparse structure on rather than setting constraints on the values of the non-zero off-diagonal elements in red: those values are updated freely during the variational optimisation (Zhao & Curtis, 2024b).
In PSVI, we can impose any desired correlation structure on by setting only the corresponding off-diagonal blocks as unknown hyperparameters and optimising them. The total number of parameters to define can be hugely reduced compared to that in full rank ADVI. The covariance matrix obtained in this way then also represents a sparse correlation structure with specific non-zero off-diagonal blocks (similar to the red elements in equation 16 but located below and above the main diagonal elements). Since a priori we expect that most of the important correlations are included in PSVI, the obtained variational distribution would capture parameter correlations of interest. Thus, inference results are significantly improved compared to those from mean-field ADVI (Zhao & Curtis, 2024b).
Variational parameters and are updated by maximising the EBLO in equation 8 or equivalently minimising the KL divergence in equations 6 and 12 using gradient based optimisation methods, and their gradients can be calculated using automatic differentiation libraries (Abadi et al., 2016; Paszke et al., 2019). The expectation terms are estimated by Monte Carlo integration with a relatively small number of samples drawn from the variational distribution, because the optimisation is performed over many iterations so that statistically the parameters will converge towards the correct solution (Kucukelbir et al., 2017).
@@ 3 Application to Full Waveform Inversion
Seismic full waveform inversion (FWI) estimates subsurface physical properties, such as seismic velocities and density, using seismic waveform data (Tarantola, 1984; Fichtner et al., 2009; Virieux & Operto, 2009). FWI is a highly nonlinear and non-unique inverse problem, and thus deterministic methods often fail to find a truly representative Earth model that generates the observed data, and to estimate reliable uncertainties in the inversion results. In recent years, researchers have started to use various Bayesian inference methods to solve probabilistic FWI problems, including Monte Carlo sampling methods (Ray et al., 2016, 2018; Visser et al., 2019; Gebraad et al., 2020; Guo et al., 2020; Kotsi et al., 2020; Zhao & Sen, 2021; Khoshkholgh et al., 2022; Biswas & Sen, 2022; Fu & Innanen, 2022; de Lima et al., 2023a, b; Berti et al., 2023) and variational inference (Zhang & Curtis, 2021b; Bates et al., 2022; Wang et al., 2023; Lomas et al., 2023; Izzatullah et al., 2023; Zhao & Curtis, 2024b; Yin et al., 2024b). Bayesian FWI often requires huge computational resources since (1) the dimensionality (number of unknown parameters) of an FWI problem is usually high (Curtis & Lomax, 2001), and (2) the forward and adjoint simulations are expensive (Wang et al., 2019; Zhao et al., 2020). Therefore, reducing computational overhead (yet still obtaining reasonably accurate inversion results) is a top priority in Bayesian FWI, especially when we have multiple types of prior information.
@@
We apply variational prior replacement (VPR) to a 2D Bayesian acoustic FWI problem to explore its effectiveness. Figure 1a displays the true velocity model used in the following tests, which is obtained by truncating and downsampling the original Marmousi model (Martin et al., 2006) to a grid size of 110 250 cells, with each cell measuring 20m in both horizontal and vertical directions. We place 12 sources (red stars in Figure 1a) on the surface with a spacing of 400m, and 250 receivers (white line in Figure 1a) on the seabed (200m depth) with a horizontal interval of 20m. The observed data contain a 4s long waveform with a sample interval of 2ms, which are generated by solving a 2D acoustic wave equation using a time-domain finite difference method. The source function is a Ricker wavelet with a dominant frequency of 10 Hz.
For Bayesian inference we define a Gaussian likelihood function to represent data uncertainties
@@
In this example, the covariance matrix of data noise is assumed to be a diagonal matrix (uncorrelated data noise). The same finite difference forward modelling method is used to calculate synthetic data , and data-model gradients are computed using the adjoint-state method (Plessix, 2006). During forward and adjoint simulations, we fix velocity values in the water layer at their true values. Prior distributions used in this study are discussed below.
@@ 3.1 Prior information
We consider three different types of prior information for the above Bayesian FWI problem. We first define a Uniform prior distribution for the velocity values at each grid cell, and its lower and upper bounds at each depth are displayed in Figure 1b, similar to that used in Zhang & Curtis (2021b). This is a non-informative (weak) and thus broad prior distribution with no correlations between neighbouring cells. It has the advantage that any type of velocity contrast between neighbouring cells would be consistent with this prior pdf (all model samples have the same prior probability values as long as they lie within the prior bounds) and hence can in principle be discriminated only by the observed waveform data.
@@
where is a normalisation constant, and is a diagonal matrix with its diagonal elements controlling the strength of the spatial smoothness. This can be interpreted as applying a Tikhonov (regularisation) matrix to (Golub et al., 1999; Aghamiry et al., 2018). Then the smoothed prior distribution can be written as
where is the Uniform distribution defined above, and is another normalisation constant which can be absorbed into the evidence term in Bayes’ rule. So we do not need to calculate its exact value.
This prior distribution embodies strong prior information in which model samples with smaller velocity contrasts between spatially neighbouring cells have higher probability values. Therefore, velocities in neighbouring cells should be positively correlated. This information may or may not be advantageous depending on the true (geological) prior information about the form of the velocity structure being estimated. Yet it effectively reduces the hypervolume of parameter space spanned by significantly non-zero values of the posterior pdf, and is more informative than the Uniform prior distribution . More detailed comparisons about these two prior distributions and their effects on the posterior pdfs can be found in Earp & Curtis (2020).
@@
To achieve this, we first select a set of realistic geological images. Figure 2 displays one such image (John, 2017), and while the images are all at much smaller scale than the Marmousi model was designed to represent, we assume scale invariance of geological correlations (only for the purposes of this test). From each of the pictures, we randomly sample 1000 subimages using a window with 20 20 pixels, which together represent a local correlation structure between parameters. Figure 3a shows the calculated full correlation matrix with a size of 400 400, each element denoting correlation information of one parameter pair within the 20 20 window. To analysis a more detailed structure of this correlation matrix, Figures 3b and 3c present its first 60 60 elements and 20 20 elements, respectively. Note that we reshape the 20 20 (2D) images into 1D vectors in a row-major order (i.e., for each training image the first 20 elements of the 1D vector comprise the first row of the 2D image, the second 20 elements comprise the second row, and so on). Therefore, off-diagonal blocks observed in Figures 3a and 3b represent correlation information in the vertical direction; only one such obvious block is visible, meaning that vertical correlations exist predominantly between vertically adjacent cells and decay rapidly with greater inter-cell distance. On the other hand, off-diagonals directly below and above the main diagonal elements (displayed in Figure 3c) denote horizontal correlations, where three or four strong off-diagonal elements have correlation values larger than 0.7, implying that strong horizontal correlations exist within approximately 4 neighbouring cells. Such horizontally smoother and vertically rougher correlation features are also clear in Figure 2.
@@
Figure 4 displays one random sample drawn from each of the three prior distributions. Since no spatial correlation is considered in the Uniform prior distribution, we observe large velocity contrasts between neighbouring cells in Figure 4a. The smoothed and geological prior distributions impose spatially correlated information, thus the prior samples presented in Figures 4b and Figures 4c are spatially smoother. In addition, local velocity structures in Figure 4c show rectangular patterns with larger sizes in the horizontal direction and smaller sizes in the vertical direction due to horizontal smoothness and vertical roughness as represented in the geological prior pdf and illustrated in Figures 2 and 3. This pattern is not observed in Figure 4b since in we impose the same magnitude of smoothness in both horizontal and vertical directions. Figure 4 thus proves that the three prior distributions encapsulate significantly different prior information that we may wish to inject into FWI inversion results.

### v2

@@ 2.4 Variational Prior Replacement
Solving the second (quasi) Bayesian problem in equation 5 requires to be evaluated. On the right hand side of equation 5, while in most cases both old and new prior probability values can be calculated efficiently, the main challenge lies in computing for a new model without invoking equation 4 (Bayes’ rule) which involves forward simulation of data corresponding to in order to evaluate the likelihood. Otherwise prior replacement would reduce to solving multiple independent Bayesian inverse problems with different prior distributions.
We use variational inference to solve the first Bayesian problem described in equation 4, providing a known and parametrised probability distribution (called the variational distribution according to the previous section) that approximates the old posterior pdf:
This distribution is found by solving a variational optimisation problem described in equation 9. Once we obtain , we can use it to replace in equation 5. Since is only an approximation (rather than exactly equal) to , we end up with an approximate expression for :
@@
This means that to estimate we do not need to evaluate and so avoid the calculation of likelihood function. Since forward simulation is the most computationally expensive component in an inverse problem, evaluating would normally be far cheaper than evaluating . Equation 11 indicates that we can estimate from , up to a normalisation constant which can be absorbed into .
@@
Note that not all variational inference methods can provide a variational distribution ( in this case) whose probability value can be evaluated easily. For example, Stein variational gradient descent (SVGD – Liu & Wang, 2016) and its stochastic version (sSVGD – Gallego & Insua, 2018) iteratively update a set of samples (also called particles) such that they become distributed according to an approximation to the posterior distribution. The output is the optimised set of particles, which are used to estimate statistical properties of the posterior distribution. However, evaluating the probability value for a particular is not at all straightforward with these methods.
@@
An alternative suite of variational methods approximates the posterior distribution as a known structure with given complexity (for example a Gaussian distribution), and can thus be expressed by a given parametric (often analytic) representation. Variational inference finds the optimal values of hyperparameters that control the parametric expression, thus defining a variational distribution that best approximates the true posterior pdf. Since we obtain a parametric (closed form) expression for the variational distribution, we can easily evaluate its probability value for any model . We refer to this kind of variational method as parametric variational inference (Sjölund, 2023). Examples of typical parametric variational inference methods include automatic differentiation variational inference (ADVI – Kucukelbir et al., 2017), normalising flows (Rezende & Mohamed, 2015), boosting variational inference (BVI – Guo et al., 2016; Miller et al., 2017), and physically structured variational inference (PSVI – Zhao & Curtis, 2024b).
@@
If is constructed using a parametric variational inference method then its probability value can be calculated easily, and equation 11 can in principle be evaluated using any probabilistic inference method since the probability value can be approximated efficiently. However, even though can be evaluated efficiently, the curse of dimensionality may nevertheless make the problem expensive, if not impossible, to solve using Monte Carlo sampling methods. We therefore introduce a second variational distribution to approximate the new posterior distribution given the new prior information . This new variational distribution can be obtained by minimising the KL-divergence between and :
@@
where the second line is obtained by substituting equation 11 into the first line. Note that is the exact distribution from equation 11, which is then approximated by using the same approximation as in equation 11 by introducing the approximate Bayesian solution . The last term is a constant and can safely be ignored when minimising KL. This optimisation problem can be solved in exactly the same way as conventional variational problem, and the result satisfies obtained by solving
@@
Using the framework of variational inference, the two main operations in the original prior replacement problem described in equations 4 and 5 are converted into two variational problems to estimate and , containing two approximate steps. We therefore call this new methodology variational prior replacement (VPR).
Note that equation 10 illustrates that the variational solution to the old Bayesian problem is an approximation, and VPR makes an additional approximation. Even if equals or if we somehow find an exact and analytic solution for , we still need to introduce to approximate the true posterior pdf by minimising the KL divergence KL in equation 12 since direct calculation of using equation 11 requires the normalisation constant to be evaluated which is intractable in high dimensional inverse problems. While the first operation (equation 10) which estimates must be performed using a parametric variational method so that its probability value can be evaluated in equations 12 and 13, the second problem can be solved using any variational method.
@@
The most expensive step in the VPR algorithm is to solve the first variational problem, because this requires the likelihood term (forward function) to be calculated. Provided that the support of all other prior pdf’s are subsets of this support, then this step needs to be performed only once, after which we can replace prior information rapidly whenever it changes. This makes the proposed method attractive in real problems, especially when multiple different prior distributions are possible for a single observed dataset (Earp & Curtis, 2020; Bloem et al., 2024).
@@ 2.5 Physically Structured Variational Inference (PSVI)
In this paper, we use physically structured variational inference (PSVI – Zhao & Curtis, 2024b) to solve the two variational problems to calculate both and . PSVI is an efficient parametric variational inference method that defines a Gaussian variational family with a physics-based correlation structure. When the model parameters to be estimated have physical constrains (for example, seismic velocity should be a positive number and earthquake source location should be below the Earth’s surface), a bijective function (an invertible transform) is usually applied to the Gaussian random variables to ensure that the transformed model parameters satisfy their physical constrains. For example, the following logit functions
@@
A Gaussian variational distribution is defined by a mean vector and a covariance matrix . To ensure that always remains positive semi-definite, we re-parametrise it using a Cholesky factorisation , where is a lower triangular matrix. A full covariance matrix can be constructed to include correlation information between pairs of model parameters. However, this incurs huge memory requirements and computational costs (for an dimensional problem, requires real-valued entries). Alternatively, a mean-field (factorised) Gaussian variational approximation may be used for high dimensional problems, which defines a diagonal covariance matrix, thus ignoring all correlations between model parameters. These two options are respectively referred to as full rank ADVI and mean-field ADVI in Kucukelbir et al. (2017). Unfortunately for the full waveform inversion (FWI) problems considered in this paper, mean-field ADVI normally underestimates uncertainties of the posterior distribution whereas full rank ADVI is intractable due to the dimensionality of models (Zhang et al., 2023; Zhao & Curtis, 2024b).
PSVI embodies a method with intermediate cost that lies between mean-field ADVI and full rank ADVI, by modelling only the most important (dominant) correlation information in model vector , guided by physical properties (prior knowledge) of imaging problems. Specifically, in spatial inverse (imaging) problems, model correlations are shown to be strong mainly between pairs of locations that are in spatial proximity to each other, and the magnitude of correlations decreases rapidly as the distance between two locations increases (Gebraad et al., 2020; Zhang & Curtis, 2021a; Biswas & Sen, 2022). This suggests that it might be sufficient to model correlations only between parameters that define properties which are spatially close (e.g., for FWI, parameters of cells that lie within a dominant wavelength of one another), and ignore correlations between those that are further apart.
@@
For each element, the first subscript indicates a block of off-diagonal elements that are rows below the main diagonal (i.e., at an offset of from the main diagonal), and the second subscript indicates that is the th element of that off-diagonal block. In equation 16, sparsely distributed off-diagonal elements in red are used to capture main correlations between parameter pairs that are assumed to be important (in this case, are spatially close), and their values are optimised during variational inference. All other off-diagonal elements of are set to be zero by assuming independence of the corresponding model parameter pairs. Note that we only impose a sparse structure on rather than setting constraints on the values of the non-zero off-diagonal elements in red: those values are updated freely during the variational optimisation (Zhao & Curtis, 2024b).
In PSVI, we can impose any desired correlation structure on by setting only the corresponding off-diagonal blocks as unknown hyperparameters and optimising them. The total number of parameters to define can thus be greatly reduced compared to that in full rank ADVI. The covariance matrix obtained in this way then also represents a sparse correlation structure with specific non-zero off-diagonal blocks (similar to the red elements in equation 16 but located below and above the main diagonal elements). Since a priori we expect that most of the important correlations are included in PSVI, the obtained variational distribution would capture parameter correlations of interest. Thus, inference results are significantly improved compared to those from mean-field ADVI (Zhao & Curtis, 2024b).
Variational parameters and are updated by maximising the EBLO in equation 8 or equivalently minimising the KL divergence in equations 6 and 12 using gradient based optimisation methods, and their gradients are calculated using automatic differentiation libraries (Abadi et al., 2016; Paszke et al., 2019). The expectation terms are estimated by Monte Carlo integration with a relatively small number of samples drawn from the variational distribution, because the optimisation is performed over many iterations so that statistically the parameters will converge towards the correct solution (Kucukelbir et al., 2017).
@@ 3 Application to Full Waveform Inversion
Seismic full waveform inversion (FWI) estimates subsurface physical properties, such as seismic velocities and density, using seismic waveform data (Tarantola, 1984; Fichtner et al., 2009; Virieux & Operto, 2009). FWI is a highly nonlinear and non-unique inverse problem, and thus deterministic methods often fail to find a truly representative Earth model that generates the observed data, and to estimate reliable uncertainties in the inversion results. In recent years, researchers have started to use various Bayesian inference methods to solve probabilistic FWI problems, including Monte Carlo sampling methods (Ray et al., 2016, 2018; Visser et al., 2019; Gebraad et al., 2020; Guo et al., 2020; Kotsi et al., 2020; Zhao & Sen, 2021; Khoshkholgh et al., 2022; Biswas & Sen, 2022; Fu & Innanen, 2022; de Lima et al., 2023a, b; Berti et al., 2023) and variational inference (Zhang & Curtis, 2021b; Bates et al., 2022; Wang et al., 2023; Lomas et al., 2023; Izzatullah et al., 2023; Zhao & Curtis, 2024b; Yin et al., 2024b). Bayesian FWI often requires huge computational resources since (1) the dimensionality (number of unknown parameters) of an FWI problem is usually high (Curtis & Lomax, 2001), and (2) the forward and adjoint simulations are expensive (Wang et al., 2019; Zhao et al., 2020). Therefore, reducing computational overhead (while still obtaining reasonably accurate inversion results) is a top priority in Bayesian FWI, especially if we have multiple prior distributions to consider.
@@
We apply variational prior replacement (VPR) to a 2D Bayesian acoustic FWI problem to explore its effectiveness. Figure 1a displays the true velocity model used in the following tests, which is obtained by truncating and downsampling the original Marmousi model (Martin et al., 2006) to a grid size of 110 250 cells, with each cell measuring 20 m in both horizontal and vertical directions. We place 12 sources (red stars in Figure 1a) on the surface with a spacing of 400 m, and 250 receivers (white line in Figure 1a) on the seabed (200 m depth) with a horizontal interval of 20 m. The waveform data are 4 s long with a sample interval of 2 ms, which are generated by solving a 2D acoustic wave equation using a time-domain finite difference method. We further add Gaussian random noise with zero mean and a standard deviation value of 0.1 ( of the average value of the maximum amplitude of each seismic trace) to the obtained waveform data, which is treated as the observed dataset in this example. The source function is a Ricker wavelet with a dominant frequency of 10 Hz.
We define a Gaussian likelihood function to represent data uncertainties
@@
In this example, the covariance matrix of data noise is assumed to be a diagonal matrix (uncorrelated data noise), and all diagonal elements are set to be 0.1 to represent the level of noise added to the synthetic waveform data. That is to say, we assume the data covariance matrix is known. The same finite difference forward modelling method is used to calculate synthetic data , and data-model gradients are computed using the adjoint-state method (Plessix, 2006). During forward and adjoint simulations, we fix velocity values in the water layer at their true values. Prior distributions used in this study are discussed below.
@@ 3.1 Prior information
We consider three different types of prior information in the FWI problem. We first define a uniform prior distribution for the velocity values at each grid cell with lower and upper bounds at each depth displayed in Figure 1b, similar to that used in Zhang & Curtis (2021b). This is a non-informative (weak) and thus broad prior distribution with no correlations between neighbouring cells. It has the advantage that any type of velocity contrast between neighbouring cells would be consistent with this prior pdf (all model samples have the same prior probability density) as long as they lie within the prior bounds, and hence can in principle be discriminated only by comparing their consistency with the observed waveform data.
@@
where is a normalisation constant, and is a diagonal matrix with its diagonal elements controlling the strength of the spatial smoothness (larger values correspond to weaker spatial smoothness). In this paper, the diagonal elements of are set to 500. This can be interpreted as applying a Tikhonov (regularisation) matrix to (Golub et al., 1999; Aghamiry et al., 2018). Then the smoothed prior distribution can be written as
where is the uniform distribution defined above, and is another normalisation constant which can be absorbed into the evidence term in Bayes’ rule so we do not need to calculate its value.
This prior distribution embodies strong prior information in which model samples with smaller velocity contrasts between spatially neighbouring cells have higher probability values. Therefore, velocities in neighbouring cells should be positively correlated. This information may or may not be advantageous depending on the true (geological) prior information about the form of the velocity structure being estimated. Compared to the uniform prior distribution, large velocity contrasts between neighbouring cells are almost excluded by this prior information as they have relatively low prior probability values. This effectively reduces the hypervolume of parameter space spanned by significantly non-zero values of the posterior pdf. In other words, it provides more information than . More detailed comparisons of these two prior distributions and their effects on the posterior pdfs can be found in Earp & Curtis (2020).
@@
To achieve this, we first select a set of realistic geological images. Figure 2 displays one such image (John, 2017), and while the images are all at much smaller scale than the Marmousi model was designed to represent, we assume scale invariance of geological correlations (only for the purposes of this test). From each of the pictures, we randomly sample 1000 subimages using a window with 20 20 pixels, which together represent a local correlation structure between parameters. Figure 3a shows the calculated full correlation matrix with a size of 400 400, each element denoting correlation information of one parameter pair within the 20 20 window. To analyse a more detailed structure of this correlation matrix, Figures 3b and 3c present its first 60 60 elements and 20 20 elements, respectively. Note that we reshape the 20 20 (2D) images into 1D vectors in a row-major order (i.e., for each training image the first 20 elements of the 1D vector comprise the first row of the 2D image, the second 20 elements comprise the second row, and so on). Therefore, off-diagonal blocks observed in Figures 3a and 3b represent correlation information in the vertical direction; only one such obvious block is visible, meaning that vertical correlations exist predominantly between vertically adjacent cells and decay rapidly with greater inter-cell distance. On the other hand, off-diagonals directly below and above the main diagonal elements (displayed in Figure 3c) denote horizontal correlations: three or four strong off-diagonal elements have correlation values larger than 0.7, implying that strong horizontal correlations exist within approximately 4 neighbouring cells. Such horizontally smoother and vertically rougher correlation features are also clear in Figure 2.
@@
Figure 4 displays one random sample drawn from each of the three prior distributions. Since no spatial correlation is considered in the uniform prior distribution, we observe large velocity contrasts between neighbouring cells in Figure 4a. The smoothed and geological prior distributions impose spatially correlated information, thus the prior samples presented in Figures 4b and Figures 4c are spatially smoother. In addition, local velocity structures in Figure 4c show rectangular patterns with larger sizes in the horizontal direction and smaller sizes in the vertical direction due to horizontal smoothness and vertical roughness as represented in the geological prior pdf and illustrated in Figures 2 and 3. This pattern is not observed in Figure 4b since in we impose the same magnitude of smoothness in both horizontal and vertical directions (this was not a requirement, but in equation 19 we used equal horizontal and vertical smoothing to contrast with the geological prior pdf). Figure 4 thus proves that the three prior distributions encapsulate significantly different prior information that we may wish to inject into FWI inversion results.